OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -4, -16, -64, -256, ...

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@iambatman @Ashleyisakitty

OpenStudy (campbell_st):

well the sequence is geometric... do you know the common ratio...?

OpenStudy (solomonzelman):

\(\large\color{black}{ a_n=a_1\times r^{n-1} }\) this is the equatioon you need to find. You just need to determine the \(\large\color{black}{ a_1 }\) and the \(\large\color{black}{ r }\) .

OpenStudy (solomonzelman):

( \(\large\color{black}{ a_1 }\) is the first term of the sequence.) ( \(\large\color{black}{ r }\) is the common ratio of the sequence. )

OpenStudy (anonymous):

Yea thats what I have in my notes but the answers dont show it with the ^ part

OpenStudy (solomonzelman):

what do you mean don't show it when the ^ part? yu mean you don't have the \(\large\color{black}{ r^{\color{red}{n-1}} }\) (in red) ?

OpenStudy (solomonzelman):

not when, it is with.

OpenStudy (anonymous):

yea there isnt that

OpenStudy (solomonzelman):

well, you either need the: \(\large\color{blue}{ a_n=a_1\times r^{n-1} }\) or \(\large\color{green}{ a_n=a_{n-1}\times r }\) .

OpenStudy (anonymous):

let me try to work it out though

OpenStudy (anonymous):

then would it be an = -4 • 4^n - 1

OpenStudy (solomonzelman):

yes, it is exactly that: Just making sure you meant what I thought you mean. (lol) \(\large\color{red}{ a_n=(-4)\times4^{n-1} }\)

OpenStudy (solomonzelman):

that is correct.

OpenStudy (anonymous):

Oh okay cool thanks :)

OpenStudy (solomonzelman):

Anytime....!

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -14, -11, -8, ... For this one can you check my answer

OpenStudy (anonymous):

an = -17 + 3(n - 1) @SolomonZelman

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yup, that's is right again:)

OpenStudy (anonymous):

but this time there is no ^ ?

OpenStudy (solomonzelman):

yes, there is no ^.

OpenStudy (anonymous):

when do I know when there will be a ^

OpenStudy (solomonzelman):

I lost connection.

OpenStudy (solomonzelman):

Okay, so in any arithmetic sequence you know that the pattern is: \(\large\color{black}{ a_1\times r=a_2 }\) \(\large\color{red}{ a_1\times r^2=a_3 }\) \(\large\color{black}{ a_1\times r^3=a_4 }\) \(\large\color{red}{ a_1\times r^4=a_3 }\) you are multiplying times the common ratio each time. that is why you have the ^.

OpenStudy (solomonzelman):

the last red row is supposed to be a5 at the end, not a3.

OpenStudy (solomonzelman):

Okay, now, an arithmetic sequence. You are adding the common difference and this is what happens. \(\large\color{black}{ a_1+ d=a_2 }\) \(\large\color{red}{ a_1+2d=a_3 }\) \(\large\color{black}{ a_1+3d=a_4 }\) \(\large\color{red}{ a_1+4d=a_5 }\) that is why you don't have the ^.

OpenStudy (anonymous):

Oh okay so depending on how the question is asked whether arithmetic or geometric

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

oh okay thankyou

OpenStudy (solomonzelman):

You welcome!